OpenStudy (anonymous):

The sum of two consecutive integers is 41. What are the integers?

OpenStudy (hoblos):

n+ (n+1) = 41 2n=40 n=20

OpenStudy (hoblos):

so they are 20 and 21

OpenStudy (anonymous):

Which let statement(s) would you use? a.) Let x = the integers. b.) Let x= first integer. Let x+1=2nd integer c.) Let x= first integer. Let x+2=2nd integer

OpenStudy (hoblos):

since they are two consecutive integers you must choose x= first integer and x+1=2nd integer

OpenStudy (anonymous):

@hoblos Thanks man appreciate you

OpenStudy (hoblos):

glad to help :D

OpenStudy (campbell_st):

a simple method is 41/2 = 20.5 the integers are either side of the average so 20 and 21